OpenStudy (anonymous):

Evaluate 48 – r for r = 12. A 36 B 60 C 72

OpenStudy (anonymous):

48-12=?

Parth (parthkohli):

\[\left.48 - r\right|_{r = 12} = 48 - 12\]